# Eraser: Jailbreaking Defense in Large Language Models via Unlearning Harmful Knowledge

### v1

@@ Abstract
Jailbreaking attacks can enable Large Language Models (LLMs) to bypass the safeguard and generate harmful content. Existing jailbreaking defense methods have failed to address the fundamental issue that harmful knowledge resides within the model, leading to potential jailbreak risks for LLMs. In this paper, we propose a novel defense method called Eraser, which mainly includes three goals: unlearning harmful knowledge, retaining general knowledge, and maintaining safety alignment. The intuition is that if an LLM forgets the specific knowledge required to answer a harmful question, it will no longer have the ability to answer harmful questions. The training of Erase does not actually require the model’s own harmful knowledge, and it can benefit from unlearning general answers related to harmful queries, which means it does not need assistance from the red team. The experimental results show that Eraser can significantly reduce the jailbreaking success rate for various attacks without compromising the general capabilities of the model.
@@ 1 Introduction
In light of this, the intuition of our method is removing the harmful knowledge from LLMs as illustrated in Figure 1 (c). We propose Eraser, a jailbreaking defense method that mainly includes three goals: unlearning harmful knowledge, retaining general knowledge, and maintaining safety alignment to harmful inquires. Specifically, we perform gradient ascent on harmful answers in a simulated jailbreaking mode, retain general knowledge by preserving the ability to understand entities, and enhance safety alignment by maintaining the ability to reject harmful questions. Experimental results have shown that the proposed method can significantly reduce the success rate of various jailbreaking attacks without compromising the performance on other tasks.
@@
Our experimental results can be reproduced at https://anonymous.4open.science/r/Eraser-537E/
@@ 2.2 LLM Unlearning
Machine unlearning methods are designed to remove specified knowledge that has been learned by a model Bourtoule et al. (2021). LLMs are trained on massive training data, re-training LLMs is obviously not a solution for forgetting specific knowledge. Using machine unlearning methods to mitigate the privacy exposure or poisoning attack on LLMs has become a promising research direction Jang et al. (2023); Chen and Yang (2023); Eldan and Russinovich (2023). Some recent work attempted to solve the harmful output problem using unlearning. Zhou et al. (2023) assumed that there were harmful instructions in the SFT dataset and attempted to make harmful behaviors unlearnable during the SFT process. The most relevant work to our work is (Yao et al., 2023), which uses unlearning to remove harmful responses, erase copyright-protected content, and eliminate hallucination from an unaligned model. Yao et al. (2023) considered the LLM unlearning as an alignment method, an alternative to RLHF In contrast, we consider the LLM unlearning as a post-hoc defense strategy against jailbreaking on an aligned model. (Yao et al., 2023) can be considered as our baseline.
@@ 3.5 Overall objective
Compared to preserving model capacity, unlearning knowledge is a much easier objective, so striking a balance among the three goals is challenging. In §4.5, we observe that prolonged unlearning training can have a detrimental effect on the model’s performance over time. Therefore, we aim to set a constraint for the unlearning objective and focus on optimizing the remaining two objectives after sufficient unlearning training:
@@ 4.1 Experimental Setup
Implementation Details. We employ Llama2-chat-7b (Touvron et al., 2023) as the base model which has undergone thorough safety alignment training. The proposed method was trained using LORA (Hu et al., 2021). During the training process, was set to 2, while and were set to 1. The batch size was fixed at 50 samples, and texts exceeding 2048 tokens were truncated. We applied the AdamW optimizer with 1e-4 learning rate and 1e-3 weight decay. The number of training epochs is set to . The checkpoint with the lowest training loss was selected for inference. For RSFT, the training settings were the same as our method. For GAM, we mostly followed the author’s settings, except for setting the maximum gradient ascent loss to 2 to accommodate the AdvBench dataset. For the attack methods AutoDAN and PAIR, we limited the maximum search steps to 20 and 25, respectively, and modified the criterion for determining whether a jailbreak has occurred to be the same as ours. That is, judging based on LLMs’ feedback.
@@ 4.2 Main Results
| ERASER | 0.50 | 1.03 | 2.88 | 1.09 | 8.26 | 1.33 | |
| AdvExtent | Base model | 23.74 | 1.86 | 44.36 | 2.65 | 17.78 | 1.65 |
| GAM (Yao et al., 2023) | 29.49 | 1.99 | 27.33 | 1.97 | 2.80 | 1.10 | |
| RSFT (Deng et al., 2023) | 0.00 | 1.00 | 2.87 | 1.09 | 2.27 | 1.09 | |
@@ 4.3 Ablation Study
Table 3 shows the experimental results. Compared to Eraser, Eraser w/o show a significant increase in ASR, indicating the effectiveness of against jailbreaking attacks. GA, which only uses gradient ascent as the goal, exhibits excellent defense performance, but its general capability is severely impaired. With the addition of the target , the general capability of Eraser w/o is mostly restored, but some ASR increase occurs due to the absence of the goal. Eraser without experiences a decrease in general performance but still outperform GA significantly, possibly due to the compensating for the model’s general language proficiency. We can further draw the following conclusions: the random prefix/suffix enhances the model’s defensive capability, compensates for the general capability, and further improves the defensive capability of the model.
@@ 4.4 What has Contributed to Defensive Capabilities?
Based on the observations above, we speculate that the sources of defensive capabilities can be diverse. Forgetting harmful text can contribute to defensive capabilities, which is a source of Eraser defense. The reason why GA w/ random brings defensive capabilities may be due to the disruption of the model’s general performance, as Eraser w/ random loses its defensive capabilities by compensating for general performance. The underlying logic is the trade-off between harmfulness and usefulness.
@@ Appendix B Compared to harmful behavior filtering method
Approaches |
Latency |
Refusal rate |
Base model |
6.73 |
0.00 |
Eraser |
6.47 |
0.00 |
RA-LLM |
11.53 |
8.00 |
@@ Appendix C The quantitative analysis of similar questions in Figure 5
Approaches |
Refusal rate |
Base model |
8.00 |
Eraser |
8.66 |
GAM |
45.63 |
RSFT |
48.99 |

### v2

@@ Abstract
Jailbreaking attacks can enable Large Language Models (LLMs) to bypass the safeguard and generate harmful content. Existing jailbreaking defense methods have failed to address the fundamental issue that harmful knowledge resides within the model, leading to potential jailbreak risks for LLMs. In this paper, we propose a novel defense method called Eraser, which mainly includes three goals: unlearning harmful knowledge, retaining general knowledge, and maintaining safety alignment. The intuition is that if an LLM forgets the specific knowledge required to answer a harmful question, it will no longer have the ability to answer harmful questions. The training of Erase does not actually require the model’s own harmful knowledge, and it can benefit from unlearning general answers related to harmful queries, which means it does not need assistance from the red team. The experimental results show that Eraser can significantly reduce the jailbreaking success rate for various attacks without compromising the general capabilities of the model. Our codes are available at https://github.com/ZeroNLP/Eraser.
@@ 1 Introduction
In light of this, the intuition of our method is removing the harmful knowledge from LLMs as illustrated in Figure 1 (c). We propose Eraser, a jailbreaking defense method that mainly includes three goals: unlearning harmful knowledge, retaining general knowledge, and maintaining safety alignment to harmful inquires. Specifically, we perform gradient ascent on harmful answers in a simulated jailbreaking mode, retain general knowledge by preserving the ability to understand entities, and enhance safety alignment by maintaining the ability to reject harmful questions.
Experimental results have shown that the proposed method can significantly reduce the success rate of various jailbreaking attacks without compromising the performance on other tasks.
@@ 2.2 LLM Unlearning
Machine unlearning methods are designed to remove specified knowledge that has been learned by a model Bourtoule et al. (2021). LLMs are trained on massive training data, re-training LLMs is obviously not a solution for forgetting specific knowledge. Using machine unlearning methods to mitigate the privacy exposure or poisoning attack on LLMs has become a promising research direction Jang et al. (2023); Chen and Yang (2023); Eldan and Russinovich (2023). Some recent work attempted to solve the harmful output problem using unlearning. Zhou et al. (2023) assumed that there were harmful instructions in the SFT dataset and attempted to make harmful behaviors unlearnable during the SFT process. The most relevant work to our work is (Yao et al., 2023), which uses unlearning to remove harmful responses, erase copyright-protected content, and eliminate hallucination from an unaligned model. However, Yao et al. considered the LLM unlearning as an alignment method, an alternative to RLHF. In contrast, we consider the LLM unlearning as a post-hoc defense strategy against jailbreaking on an aligned model.
@@ 3.5 Overall objective
Compared to preserving model capability, unlearning knowledge is a much easier objective, so striking a balance among the three goals is challenging.
In §4.5, we observe that prolonged unlearning training can have a detrimental effect on the model’s performance over time.
Therefore, we aim to set a constraint for the unlearning objective and focus on optimizing the remaining two objectives after sufficient unlearning training:
@@ 4.1 Experimental Setup
Implementation Details. We employ Llama2-chat-7b Touvron et al. (2023) as the base model which has undergone thorough safety alignment training. The proposed method was trained using LORA Hu et al. (2021). During the training process, was set to 2, the batch size was fixed at 64 samples, and texts exceeding 2048 tokens were truncated. We applied the AdamW optimizer with 2e-5 learning rate. The number of training epochs is set to 5. The checkpoint with the lowest training loss was selected for inference. For RSFT, we employ a learning rate of 1e-4 and a weight decay of 1e-3. For GAM, we mostly followed the author’s settings, except for stopping the training when the gradient reaches 2 to accommodate the AdvBench dataset. For the attack methods AutoDAN, we limited the maximum search steps to 20, and modified the criterion for determining whether a jailbreak has occurred to be the same as ours. That is, judging based on LLMs’ feedback.
@@ 4.2 Main Results
| ERASER | 0.57 | 1.03 | 2.88 | 1.09 | 8.26 | 1.33 | |
| AdvExtent | Base model | 23.74 | 1.86 | 44.36 | 2.65 | 17.74 | 1.65 |
| GAM (Yao et al., 2023) | 29.49 | 1.99 | 27.33 | 1.97 | 2.87 | 1.11 | |
| RSFT (Deng et al., 2023) | 0.00 | 1.00 | 2.87 | 1.09 | 2.15 | 1.09 | |
@@ 4.3 Ablation Study
Table 3 shows the experimental results. Compared to Eraser, Eraser w/o show a significant increase in ASR, indicating the effectiveness of against jailbreaking attacks. GA, which only uses gradient ascent as the goal, exhibits excellent defense performance, but its general capability is severely impaired. With the addition of the target , the general capability of Eraser w/o is mostly restored, but some ASR increase occurs due to the absence of the goal. Eraser w/o experiences a decrease in general performance but still outperform GA significantly, possibly due to the compensating for the model’s general language proficiency. We can further draw the following conclusions: the random prefix/suffix enhances the model’s defensive capability, compensates for the general capability, and further improves the defensive capability of the model.
@@ 4.4 What has Contributed to Defensive Capabilities?
Based on the observations above, we speculate that the sources of defensive capabilities can be diverse. Forgetting harmful text can contribute to defensive capabilities, which is a source of Eraser defense. The reason why GA w/ random brings defensive capabilities may be due to the disruption of the model’s general performance, as Eraser w/ random loses its defensive capabilities by compensating for general performance. The underlying logic is the trade-off between harmfulness and usefulness. The model loses the ability to follow instructions, naturally losing the ability to follow harmful instructions as well.
Considering that GA reduces the general ability by 1.94% while decreasing the ASR of AIM attacks from 19.61% to 5.38%, and its implementation cost is extremely low, requiring only the random generation of some data to unlearning, defensive capability appears to be a relatively easily acquired attribute. Recall that RSFT’s 2.28% reduction in general capability, its good defense performance is not surprising. In comparison, Eraser’s ability to maintain general capability is particularly valuable.
@@ Appendix B Compared to harmful behavior filtering method
| Approaches | Latency | Refusal rate |
| Base model | 6.73 | 0.00 |
| Eraser | 6.47 | 0.00 |
| RA-LLM | 11.53 | 8.00 |
@@ Appendix C The quantitative analysis of similar questions in Figure 5
| Approaches | Refusal rate |
| Base model | 8.00 |
| Eraser | 8.66 |
| GAM | 45.63 |
| RSFT | 48.99 |